OpenStudy (anonymous):

Ive been stuck all morning... Which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring? -b b2 - 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x2 + 7x + 3 = 0

OpenStudy (anonymous):

the first answer is b2-4ac the second answer is b^2-4ac=(7)^2-4(2)(3)=49-24=25 does that clear it up for you at all? or do u need more help?